# Syzygies and Koszul modules in geometry

### v1

@@ Acknowledgments
This paper is dedicated to the memory of Ş. Papadima who first introduced me to Koszul modules. Without his insight the connections between syzygies and Koszul modules would probably not have been discovered. I have greatly profited from discussions with M. Aprodu, M. Kemeny, R. Lazarsfeld, J. Park, C. Raicu, F.-O. Schreyer, A. Suciu, C. Voisin and J. Weyman related to this circle of ideas. I am particularly grateful to C. Raicu for many useful comments on the first version of this paper. This work was supported by the Berlin Mathematics Research Center MATH+ and by the ERC Advanced Grant SYZYGY.
@@ 4.2. Suciu’s conjecture for hyperplane arrangement groups
and set $\mathcal{P}_{\mathcal{M}}:=\langle u_{2}-u_{1},\ldots,u_{k}-u_{1}\rangle\subseteq H^{1}(M(\mathcal{A}),\mathbb{C})$. It is shown in [FY, Theorem 2.4] that $\mathcal{P}_{\mathcal{M}}$ is an isotropic (k1)(k-1)-dimensional component of the resonance $\mathcal{R}\bigl(\mathcal{A})\bigr)$. Furthermore, if $\mathcal{B}$ is a subarrangement of $\mathcal{A}$ and $\mathcal{N}$ is a multinet of $\mathcal{B}$, then $\mathcal{P}_{\mathcal{N}}\subseteq H^{1}(M(\mathcal{A}),\mathbb{C})$ is still a component of $\mathcal{R}(\mathcal{A})$. In fact all the components of the resonance appear in this way and one has the following set-theoretic equality
@@
Suciu [Su1] conjectured that the Chen rank formula (17) holds for every hyperplane arrangement $\mathcal{A}$ for $q\gg 0$. Theorem 4.11 is an optimal effective version of Suciu’s Conjecture for arrangements having strongly isotropic resonance.
As pointed out in [AFRS2], Theorem 4.11 can be applied to all graphic arrangements. Precisely, to a simple graph $\Gamma=(V,E)$ on a vertex set $V=\{1,\ldots,m\}$, we define the arrangement $\mathcal{A}_{\Gamma}$ consisting of those hyperplanes $H_{ij}:\{z_{i}=z_{j}\}$ in $\mathbb{C}^{m}$, when $\{i,j\}\in E$. Denoting by $\kappa_{r}$ the set of (complete) $K_{r}$-subgraphs of $\Gamma$, then it is shown in [AFRS2, Corollary 9.15] that the Chen rank of group $G(\mathcal{A}_{\Gamma})$ are given by the formula
@@
Via the interpretation (5), we have that $\theta_{q}\bigl(G(\mathcal{A})\bigr)=\mbox{dim}_{\mathbb{C}}\mbox{Tor}^{E}_{q-1}\bigl(A(\mathcal{A}),\mathbb{C}\bigr)_{-q}$. One can ask whether there is an explicit combinatorial formula, similar to Suciu’s Conjecture (Theorem 4.11) for the LCS ranks $\phi_{q}\bigl(G(\mathcal{A})\bigr)$. It has been shown by Falk and Randell [FR] that one has a formula
(18) $\prod_{q=1}^{\infty}(1-t^{q})^{\phi_{q}(G(\mathcal{A}))}=\sum_{i=1}^{\infty}b_{i}\cdot t^{i},$
where $b_{i}:=\mbox{dim }\mbox{Tor }_{i}^{A(\mathcal{A})}\bigl(\mathbb{C},\mathbb{C})_{-i}$, if and only if $A(\mathcal{A})$ is a Koszul algebra. This covers several important cases, but Peeva [Pe] showed that the formula (18) could not possibly hold for all arrangements $\mathcal{A}$. This invites the question whether one can find a combinatorial formula for the ranks $\phi_{q}(\mathcal{A})$, at least for $q\gg 0$, in the same spirit as Suciu’s Chen Ranks Conjecture?
@@ 5.1. Green’s Conjecture
Green’s Conjecture remains wide open for arbitrary smooth curves. Voisin [V1], [V2] established Green’s Conjecture for general curves of every genus (in characteristic zero), using in an essential way the geometry of $K3$ surfaces. Her work has been essential for subsequent developments like Aprodu’s Brill-Noether-theoretic sufficient condition for Green’s Conjecture [Ap], the proof that all curves on $K3$ surfaces verify Green’s Conjecture [AF1], or the work on the Prym–Green Conjecture [FK1], [FK2], [FK3]. It is important to point out that for curves of maximal Clifford index $\mbox{Cliff}(X)=\lfloor\frac{g-1}{2}\rfloor$, the vanishing (21) predicted by Grren’s Conjecture is enough to determine entirely the Betti table of $X$, that is, the values of the Betti numbers $b_{p,q}(X,\omega_{X})$ for all $p$ and $q$. This is because via the description (19) of the Koszul cohomology groups $K_{p,q}(X,\omega_{X})$, the difference between the Betti diagrams on each diagonal are fixed, precisely
@@ 5.2. Green’s Conjecture via the tangent developable
(23) $K_{\lfloor\frac{g-1}{2}\rfloor,2}(X,\omega_{X})=0.$
@@
The linear section of $\mathcal{T}$ are $g$-cuspidal rational curves $X:=\mathcal{T}\cap H\subseteq H\cong{\textbf{P}}^{g-1}$ and the $g$-cusps correspond to the points of intersection of the hyperplane $H$ with the rational normal curve $\Gamma$. It follows from the fundamental work of Eisenbud and Harris [EH] on limit linear series that all $g$-cuspidal rational curves are general from the point of Brill–Noether theory, in particular any of their smoothing will be a curve of genus $g$ of maximal Clifford index $\lfloor\frac{g-1}{2}\rfloor$. This fact, which remarkably, holds irrespective of the position of the $g$ cusps was one of the reasons which made the tangential developable particularly attractive for trying to attack the Generic Green’s Conjecture. Furthermore, since Koszul cohomology obeys a form of the Lefschetz Hyperplane Principle [Gr, Theorem 3.7.b], one has isomorphism
@@
Schreyer observed computationally that Green’s Conjecture fails for every curve in low characteristic, the first counterexamples being for $g=7$ and $\mbox{char}(\mathbf{k})=2$, respectively for $g=9$ and $\mbox{char}(\mathbf{k})=3$. Eisenbud and Schreyer [ES] conjectured that the Generic Green’s Conjecture should hold whenever $\mbox{char}(\mathbf{k})\geq\frac{g-1}{2}$ and Theorem 5.4 not only provides an alternative proof of the Generic Green’s Conjecture in characteristic zero, but also (essentially) answers the Eisenbud–Schreyer Conjecture.
@@
Can one determine all characteristics $p=\mbox{char}(k)<\frac{g-1}{2}$, for which Green’s Conjecture fails for every smooth curve $X$ of genus $g$? Is there a connection between this question and the results in [RV, Theorem 7.1], where one establishes the surprising non-vanishing result $W_{2n-8}(V,K)\cong\mathbf{k}$, for any $\mathbf{k}$-vector space $V$ of dimension $n=3+p^{a}$, with $p=\mbox{char}(\mathbf{k})$, and for every (2n3)(2n-3)-dimensional subspace $K\subseteq\bigwedge^{2}V$ such that $\mathcal{R}(V,K)=0$? Note that this connection, even if it exists, cannot be provident by the tangential surface $\mathcal{T}$ used in [AFPRW1].
@@
is a divisor in $\mathcal{M}_{2i+3}$. This is stark contrast to Voisin’s original approach using curves on (smooth) K3 surfaces, when the even and the odd genus case required two largely independent papers [V1] and [V2].
@@
The Prym–Green Conjecture [CEFS] predicts the shape of the resolution of a general line bundle $L\in\mbox{Pic}^{2g-2}(X)$ on a general curve $X$ of genus $g$. The conjecture remains unknown for even genus, in which case it reduces to one single vanishing statement
$K_{\frac{g}{2}-2,1}(X,L)=0,\ \mbox{for a general }L\in\mathrm{Pic}^{2g-2}(X).$
The conjecture has been completely solved for odd genus, see [FK2], [FK3].
By analogy with [AFPRW1] we can ask whether there is a degeneration of such an embedded curve $X\stackrel{{\scriptstyle|L|}}{{\hookrightarrow}}{\textbf{P}}^{g-2}$, such that for this degenerate curve the Prym–Green Conjecture turns into a statement on a particular Koszul module. What should this Koszul module be?
@@ 6. The Secant Conjecture, the Gonality Conjecture and related topics
Another case recently treated in [F2] concerns the case when condition appearing in the left hand side of (34) describes a divisor in the Jacobian variety $\mbox{Pic}^{d-2g+2}(X)$, which happens precisely when $d=g+2p+3$.
@@
Let $X$ be a smooth curve of genus $g$, and fix $\left\lceil\frac{g-3}{2}\right\rceil\leq p\leq g-3$. Assume

### v2

# arXiv:2602.22493v2[math.AG] 28 Feb 2026

@@ Acknowledgments
This paper is dedicated to the memory of Ş. Papadima who first introduced me to Koszul modules. Without his insight the connections between syzygies and Koszul modules would probably not have been discovered. I have greatly profited from discussions with M. Aprodu, M. Kemeny, R. Lazarsfeld, J. Park, C. Raicu, F.-O. Schreyer, A. Suciu, C. Voisin and J. Weyman related to this circle of ideas. I am particularly grateful to M. Aprodu and C. Raicu for many useful comments on the first version of this paper. This work was supported by the Berlin Mathematics Research Center MATH+ and by the ERC Advanced Grant SYZYGY.
@@ 4.2. Suciu’s conjecture for hyperplane arrangement groups
and set $\mathcal{P}_{\mathcal{M}}:=\langle u_{2}-u_{1},\ldots,u_{k}-u_{1}\rangle\subseteq H^{1}(M(\mathcal{A}),\mathbb{C})$. It is shown in [FY, Theorem 2.4] that $\mathcal{P}_{\mathcal{M}}$ is an isotropic (k1)(k-1)-dimensional component of the resonance $\mathcal{R}(\mathcal{A})$. Furthermore, if $\mathcal{B}$ is a subarrangement of $\mathcal{A}$ and $\mathcal{N}$ is a multinet of $\mathcal{B}$, then $\mathcal{P}_{\mathcal{N}}\subseteq H^{1}(M(\mathcal{A}),\mathbb{C})$ is still a component of $\mathcal{R}(\mathcal{A})$. In fact all the components of the resonance appear in this way and one has the following set-theoretic equality
@@
Suciu [Su1] conjectured that the Chen rank formula (17) holds for every hyperplane arrangement $\mathcal{A}$ as long as $q\gg 0$. Theorem 4.11 is an optimal effective version of Suciu’s Conjecture for arrangements having strongly isotropic resonance.
As pointed out in [AFRS2], Theorem 4.11 can be applied to all graphic arrangements. Precisely, to a simple graph $\Gamma=(V,E)$ on a vertex set $V=\{1,\ldots,m\}$, we define the arrangement $\mathcal{A}_{\Gamma}$ consisting of those hyperplanes $H_{ij}:\{z_{i}=z_{j}\}$ in $\mathbb{C}^{m}$, when $\{i,j\}\in E$. Denoting by $\kappa_{r}$ the set of (complete) $K_{r}$-subgraphs of $\Gamma$, then it is shown in [AFRS2, Corollary 9.15] that the Chen rank of group $G(\mathcal{A}_{\Gamma})$ are given by the following formula, see also [SS2]:
@@
Via the interpretation (5), we have that $\theta_{q}\bigl(G(\mathcal{A})\bigr)=\mbox{dim}_{\mathbb{C}}\mbox{Tor}^{E}_{q-1}\bigl(A(\mathcal{A}),\mathbb{C}\bigr)_{-q}$. One can ask whether there is an explicit combinatorial formula, similar to Suciu’s Conjecture (Theorem 4.11) for the lower central series ranks $\phi_{q}\bigl(G(\mathcal{A})\bigr)$. It has been shown by Falk and Randell [FR], see also [SS1, 1.3] that one has a formula
(18) $\prod_{q=1}^{\infty}(1-t^{q})^{\phi_{q}(G(\mathcal{A}))}=\sum_{i=1}^{\infty}(-1)^{i}b_{i}\cdot t^{i},$
where $b_{i}:=\mbox{dim }\mbox{Tor }_{i}^{A(\mathcal{A})}\bigl(\mathbb{C},\mathbb{C})_{-i}$, if and only if $A(\mathcal{A})$ is a Koszul algebra. This covers several important cases, see again [SS1] and references therein, but Peeva [Pe] showed that the formula (18) could not possibly hold for all arrangements $\mathcal{A}$. This invites the question whether one can find a combinatorial formula for the ranks $\phi_{q}(\mathcal{A})$, at least for $q\gg 0$, in the same spirit as Suciu’s Chen Ranks Conjecture?
@@ 5.1. Green’s Conjecture
Green’s Conjecture remains wide open for arbitrary smooth curves. Voisin [V1], [V2] established Green’s Conjecture for general curves of every genus (in characteristic zero), using in an essential way the geometry of $K3$ surfaces. Her work has been essential for subsequent developments like Aprodu’s Brill-Noether-theoretic sufficient condition for Green’s Conjecture [Ap], the proof that all curves on $K3$ surfaces verify Green’s Conjecture [AF1], or the work on the Prym–Green Conjecture [FK1], [FK2], [FK3]. It is important to point out that for curves of maximal Clifford index $\mbox{Cliff}(X)=\lfloor\frac{g-1}{2}\rfloor$, the vanishing (21) predicted by Green’s Conjecture is enough to determine entirely the Betti table of $X$, that is, the values of the Betti numbers $b_{p,q}(X,\omega_{X})$ for all $p$ and $q$. This is because via the description (19) of the Koszul cohomology groups $K_{p,q}(X,\omega_{X})$, the difference between the Betti diagrams on each diagonal are fixed, precisely
@@ 5.2. Green’s Conjecture via the tangent developable
(23) $K_{\lfloor\frac{g-3}{2}\rfloor,2}(X,\omega_{X})=0.$
@@
The linear section of $\mathcal{T}$ are $g$-cuspidal rational curves $X:=\mathcal{T}\cap H\subseteq H\cong{\textbf{P}}^{g-1}$ and the $g$-cusps correspond to the points of intersection of the hyperplane $H$ with the rational normal curve $\Gamma$. It follows from the fundamental work of Eisenbud and Harris [EH] on limit linear series that all $g$-cuspidal rational curves are general from the point of Brill–Noether theory, in particular any of their smoothing will be a curve of genus $g$ of maximal Clifford index $\lfloor\frac{g-1}{2}\rfloor$. This fact, which remarkably, holds irrespective of the position of the $g$ cusps, was one of the reasons which made the tangential developable particularly attractive for trying to attack the Generic Green’s Conjecture. Furthermore, since Koszul cohomology obeys a form of the Lefschetz Hyperplane Principle [Gr, Theorem 3.7.b], one has isomorphism
@@
Schreyer observed computationally that Green’s Conjecture fails for every curve in low characteristic for infinitely many genera. The first counterexamples are $g=7$ and $\mbox{char}(\mathbf{k})=2$, respectively $g=9$ and $\mbox{char}(\mathbf{k})=3$. Eisenbud and Schreyer [ES] conjectured that the Generic Green’s Conjecture should hold whenever $\mbox{char}(\mathbf{k})\geq\frac{g-1}{2}$ and Theorem 5.4 not only provides an alternative proof of the Generic Green’s Conjecture in characteristic zero, but also (essentially) answers the Eisenbud–Schreyer Conjecture.
@@
Can one determine all characteristics $p=\mbox{char}(\mathbf{k})<\frac{g-1}{2}$, for which Green’s Conjecture fails for every smooth curve $X$ of genus $g$? Is there a connection between this question and the results in [RV, Theorem 7.1], where one establishes the surprising non-vanishing result $W_{2n-8}(V,K)\cong\mathbf{k}$, for any $\mathbf{k}$-vector space $V$ of dimension $n=3+p^{a}$, with $p=\mbox{char}(\mathbf{k})$, and for every (2n3)(2n-3)-dimensional subspace $K\subseteq\bigwedge^{2}V$ such that $\mathcal{R}(V,K)=0$? Note that this connection, if it exists, cannot be provided by the tangential surface $\mathcal{T}$ used in [AFPRW1].
@@
is a divisor in $\mathcal{M}_{2i+3}$. This is in stark contrast to Voisin’s original approach using curves on (smooth) K3 surfaces, when the even and the odd genus case required two largely independent papers [V1] and [V2].
@@
The Generic Prym–Green Conjecture [CEFS] predicts the shape of the resolution of a general line bundle $L\in\mbox{Pic}^{2g-2}(X)$ on a general curve $X$ of genus $g$. The conjecture remains unknown for even genus, in which case it reduces to one single vanishing statement
$K_{\frac{g}{2}-2,1}(X,L)=0,\ \mbox{for a general line bundle }L\in\mathrm{Pic}^{2g-2}(X).$
The original Prym–Green Conjecture has been formulated in [CEFS] for line bundles $L=\omega_{X}\otimes\eta$, where $\eta\in\mbox{Pic}^{0}(X)[2]$ is an $\ell$-torsion point in the Jacobian of $X$. For the surprising failure of this conjecture when $g=8$ and $\ell=2$, we refer to [CFVV]. The conjecture has been completely solved for odd genus, see [FK2], [FK3].
By analogy with [AFPRW1] we can ask whether there is a degeneration o an embedded curve
$X\stackrel{{\scriptstyle|L|}}{{\hookrightarrow}}{\textbf{P}}^{g-2}$
of degree 2g22g-2, such that for this degenerate curve the Prym–Green Conjecture turns into a statement on a particular Koszul module. What should this Koszul module be?
@@ 6. The Secant Conjecture, the Gonality Conjecture and related topics
Another situation recently treated in [F2] concerns the case when the condition appearing in the left hand side of (34) describes a divisor in the Jacobian variety $\mbox{Pic}^{d-2g+2}(X)$, which happens precisely when $d=g+2p+3$.
@@
Let $X$ be a smooth curve of genus $g$, and fix $\lceil\frac{g-3}{2}\rceil\leq p\leq g-3$. Assume